# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ 3.3. Unknown variance
Note that the scale of the random component in Theorem 4.6 is much larger than the error term in (3.8). As a result, our confidence intervals are much wider than this error term and normalizing the data by $\widehat{\sigma}$ does not change the validity of the confidence intervals constructed in the previous section.
@@
The Marchenko–Pastur law with explicit estimates for the reminders (see e.g., Bourgade et al. (2022)), combined with the interlacing inequalities between the eigenvalues of spiked and unspiked models, as in Corollaries 9.8 and 9.12, can be used to show that
@@ 3.4. Implications and interpretations
Confidence intervals serve two important roles in the analysis of signal strength. First, they provide a measure of uncertainty: the narrower the confidence interval, the more precisely the signal strength is estimated.
Second, confidence intervals help assess the informativeness of estimated signals. If the lower bound of a confidence interval starts at $\infty$- ∞ or at the minimal admissible value of $\theta$, (which is $\sigma^{2}$ for the spiked covariance and zero for other models), then the signal may be spurious and could reflect pure noise – that is, a situation in which no true signal is present. Alternatively, if the confidence interval is bounded away from the minimal admissible value, but contains the identification threshold $\theta^{c}$, then we know that a signal exists, but we cannot reject the null hypothesis that its strength falls below the identification cutoff. When $\theta\leq\theta^{c}$ the sample estimates of the signal directions $\mathbf{u}$ and $\mathbf{v}$ are asymptotically orthogonal to their true population counterparts (see, e.g., Paul (2007); Onatski (2012); Benaych-Georges and Nadakuditi (2012); Johnstone and Paul (2018); Bykhovskaya and Gorin (2025)), rendering the signal effectively non-informative.
@@ 4. Asymptotics through the Airy–Green function
The new asymptotics, which improves upon the Gaussian approximations (2.3), (2.9), (2.13), and (2.17), is based on a novel stochastic object we call the Airy–Green function. Its definition, along with the transition process $\mathcal{T}(\Theta)$ constructed from it, is presented in Section 4.1. Further discussion of its nature is provided in Section 4.2. Theorem on the convergence of the eigenvalue distributions in four signal plus noise models towards this object is stated in Section 4.3.
@@ 4.1. Definition of $\mathcal{G}(w)$ and $\mathcal{T}(\Theta)$
Let $Y_{N}$ be an $N\times N$ matrix of i.i.d. $\mathcal{N}(0,\tfrac{2}{N})$ Gaussian random variables and let $\lambda_{1;N}\geq\lambda_{2;N}\geq\dots\geq\lambda_{N;N}$ be the eigenvalues of $\mathbf{B}=\frac{1}{2}\left(Y_{N}+Y_{N}^{\mathsf{T}}\right)$. Then in the sense of convergence of finite-dimensional distributions
@@
Similar asymptotic results hold for the other models we consider, see further below. All existing formulae for the finite-dimensional distributions of $\{\mathfrak{a}_{i}\}_{i=1}^{\infty}$ are quite complicated and do not provide explicit distribution function, see, e.g., Forrester (2010). Nevertheless, the distribution can be sampled and tabulated, see Bornemann (2009); Bykhovskaya et al. (2024). In particular, Table 2 lists quantiles of the Tracy–Widom distribution, which describes the law of $\mathfrak{a}_{1}$.
@@
(4.2) $\mathcal{G}(w)=\lim_{x\to-\infty}\left[\left(\sum_{j:\,\mathfrak{a}_{j}>x}\frac{\xi_{j}^{2}}{w-\mathfrak{a}_{j}}\right)-\frac{2}{\pi}\sqrt{-x}\right]\leavevmode\nobreak\ ,$
@@
Note that any fixed $w\in\mathbb{C}$ is almost surely not in $\{\mathfrak{a}_{j}\}$, hence, for such $w$ the convergence holds almost surely. Eq. (4.2) and Proposition 8.3 imply that $\mathcal{G}(w)$ changes monotonically from +$\infty$+ ∞ to $\infty$- ∞ over the interval $[\mathfrak{a}_{1},+\infty)$, allowing us to state the following key definition.
@@
Figure 2 shows the simulated quantiles for the random variables $\mathcal{T}(\Theta)$ as functions of $\Theta$, or equivalently, confidence intervals for $\Theta$ as a function of $\mathcal{T}$. The underlying data is given in Table 1. These results are based on $MC=10^{6}$ Monte Carlo simulations of the $\sqrt{N}\times\sqrt{N}$ top-left corners of $N\times N$ tridiagonal matrices of Dumitriu and Edelman (2002), with a perturbed $(1,1)$ matrix element and $N=10^{8}$; see Edelman and Persson (2005, Section 1.1) and Johnstone et al. (2021, Lemma 5.2) for justifications of this approach. The figure shows that the Gaussian approximation from Proposition 4.3 performs very well for large $\Theta$, but deteriorates near $\Theta=0$.
@@ 4.2. Discussion of the definition
Let us clarify the terminology. The term “Airy” in the name $\mathcal{G}(w)$ refers to the Airy point process, whose points $\mathfrak{a}_{i}$ appear in its definition. The term “Green” stands from the tradition in random matrix theory to refer to matrix elements of the resolvent $(zI-D)^{-1}$ of a symmetric matrix $D$ as the Green’s function. Via eigenvalue decomposition, the $(1,1)$ matrix element of $(zI-D)^{-1}$ is $\sum_{i}\frac{u_{1i}^{2}}{z-d_{i}}$, where $u_{1i}$ is the first coordinate of the $i$th normalized eigenvector of $D$ corresponding to the eigenvalue $d_{i}$, making it reminiscent of the sum in (4.2).
The term “transition” in the name of $\mathcal{T}(\Theta)$ refers to its role in capturing the transition between subcritical $\theta<\theta^{c}$ and supercritical $\theta>\theta^{c}$ behavior in (2.3), (2.9), (2.13), (2.17). This phenomenon is commonly known as the BBP phase transition, following Baik et al. (2005).
@@ 4.3. Universal asymptotics for spiked models
[…]
@@
(4.8) $\kappa_{1}=\frac{1}{2}\left[{\frac{[V^{\prime}(\theta^{c})]^{2}}{\lambda^{\prime\prime}(\theta^{c})}}\right]^{\frac{1}{3}},\qquad\kappa_{2}=\left[{\frac{[\lambda^{\prime\prime}(\theta^{c})]^{2}}{V^{\prime}(\theta^{c})}}\right]^{\frac{1}{3}}\leavevmode\nobreak\ .$
@@
While the distributional limit of $\lambda_{q+1},\dots,\lambda_{r}$ can be also computed, it is of no use for the confidence intervals: the limiting random variables would depend on $\tilde{\theta}$, but not on $\theta_{q+1},\dots,\theta_{r}$.
@@
On the other hand, if one sets $\theta_{i}=\theta^{c}+\varepsilon$, then using (4.6) and Taylor expanding (noting $V(\theta^{c})=\lambda^{\prime}(\theta^{c})=0$), we expect
@@
which is readily seen to be equivalent to (4.6). When $\theta_{i}$ is close to $\theta^{c}$, $\theta_{i}=\theta^{c}+N^{-1/3}\tilde{\theta}$, Taylor expanding $\lambda(\cdot)$ and $V(\cdot)$ near $\theta^{c}$, (4.9) turns into
@@
which is the same as (4.7).
Note that $\kappa_{3}/N$ is $o(N^{-2/3})$, and therefore the choice of $\kappa_{3}$ does not affect the validity of the asymptotic formulas (4.9) and (4.10). These terms are introduced to further improve the performance of the formulas for intermediate values of $N$, cf. Johnstone (2008); Ma (2012); Johnstone and Ma (2012), which emphasize the importance of $1/N$ corrections for the practical applicability of limit theorems. The reasoning behind our choice of $\kappa_{3}$ is as follows. First, we require continuity at $\theta^{c}$; hence, (4.9) and (4.10) use the same $\kappa_{3}/N$. Second, we leverage additional information available at $q=r=1$ and $\tilde{\theta}=-N^{1/3}\theta^{c}$ for the spiked Wigner, factor, and CCA models. (For the spiked covariance model, one instead takes $\tilde{\theta}=-N^{1/3}\gamma$ and adjusts the formula accordingly.) On one hand, combining (4.10) with the asymptotic approximation (4.4), we obtain
@@
On the other hand, $\tilde{\theta}=-N^{1/3}\theta^{c}$ corresponds to $\theta=0$ (or $\theta=1$ for the spiked covariance model with $\tilde{\theta}=-N^{1/3}\gamma$), meaning that in all four models of interest, we are in the unspiked regime without a signal component. In this setting, the convergence of $\lambda_{1}$ to the Tracy–Widom distribution $\mathfrak{a}_{1}$ is well established, and $1/N$-order asymptotic corrections have been studied in Johnstone (2008); Ma (2012); Johnstone and Ma (2012). From these works, one can extract
(4.12) $\lambda_{1}\approx\lambda_{+}+\frac{\kappa_{1}}{N^{2/3}}\mathfrak{a}_{1}-\frac{1}{2N}\times\begin{cases}1&\text{ for spiked Wigner},\\ \gamma(1+\gamma)^{2}&\text{ for spiked covariance and factors},\\ \frac{(\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}-1)^{2}(\sqrt{\tau_{N}-1}+\sqrt{\tau_{M}-1})^{2}}{\tau_{N}^{2}\tau_{M}\sqrt{\tau_{N}-1}\sqrt{\tau_{M}-1}}&\text{ for CCA.}\end{cases}$
@@
We expect that (4.9) also remains asymptotically valid on all mesoscropic scales, i.e., when $\theta_{i}=\theta^{c}+N^{-\alpha}\tilde{\theta}$, $30<\alpha<1/$ 3. We omit a detailed proof.
The proof of Theorem 4.6 is given in Section 9. It begins with a rank-one perturbation equation, which expresses the eigenvalues in a signal plus noise model with $r$ spikes (the “target model”) as solutions to an algebraic equation involving a simpler model with $r-1$ spikes (the “base model”). Analyzing the asymptotic behavior of this equation leads to the following conclusion, stated informally below:
@@
We expect that our methods can be extended to handle the case of multiple ($k>1$) critical spikes, as well as the remaining largest eigenvalues $\lambda_{q+1},\lambda_{q+2},\dots$. The limiting behavior should be described by a higher-rank Airy point process, which we define recursively. The rank 00 process is the classical Airy point process $\{\mathfrak{a}_{j}\}$. The rank $1$ process $\{\mathfrak{a}_{j}^{(\Theta)}\}$ consists of all real solutions to the equation $\mathcal{G}(w)=-\Theta$; in particular, the largest point $\mathfrak{a}_{1}^{(\Theta)}$ coincides with $\mathcal{T}(\Theta)$ from Definition 4.2. We then iterate this construction: given the rank $k$ point process $\{\mathfrak{a}_{j}^{(\Theta_{1},\dots,\Theta_{k})}\}$ depending on $k$ real parameters $\Theta_{1},\dots,\Theta_{k}$, we define the rank $(k+1)$ process $\{\mathfrak{a}_{j}^{(\Theta_{1},\dots,\Theta_{k},\Theta_{k+1})}\}$ as the set of all real solutions to the equation
@@ 5. Empirical illustrations
In this section, we present three examples that illustrate the application of the procedure described in Section 3 to empirical data sets. In each case, the first step reveals a strong agreement between the histogram of eigenvalues and the corresponding theoretical curve—specifically, the Marchenko-Pastur law in the first two examples (factor models) and the Wachter law in the third (CCA). This suggests that the data aligns well with our modeling assumptions.
@@ 5.1. Industrial Production
Industrial production (IP) accounts for more than $10\%$ of the United States’ Gross Domestic Product (GDP), making it a significant component of total U.S. output. In this subsection we use data from Andreou et al. (2019), which investigates whether IP constitutes a dominant factor in U.S. economic activity. The data set contains quarterly IP growth rates across 117 sectors, spanning the period from $11977:Q$ 1 to $42011:Q$ 4. We de-mean the data and standardize each sector to have unit sample variance, thereby working with the sample correlation matrix of IP.
Figure 3 shows all eigenvalues of the standardized IP and highlights four of them (3.75, 4.26, 6.12, 30.053.754.266.1230.053.75,\,4.26,\,6.12,\,30.053.75 , 4.26 , 6.12 , 30.05) that lie to the right of the theoretical Marchenko–Pastur upper edge, $\lambda_{+}=3.68$. The largest eigenvalue, $30.05$, stands out markedly and represents a strong “market” factor. The two smallest among the four, being close to the cutoff, may reflect spurious signals arising from noise. To assess their significance, we construct $95\%$ confidence intervals. Notably, the interval for the eigenvalue at $3.86$ intersects the critical identification threshold $\theta^{c}=0.92$, indicating that we cannot reject the null hypothesis that it represents noise (or non-informative signal). For the remaining eigenvalues, the null is rejected. We thus conclude that the IP growth rate is driven by three factors: strong, semi-strong, and weak.

### v2

@@ 3.4. Unknown $\sigma^{2}$
The scale of the random component in Theorem 4.6 is much larger than the error term in (3.8). Henc3, our confidence intervals are much wider than this error term and normalizing the data by $\widehat{\sigma}$ does not change the validity of the confidence intervals of Sections 3.2, 3.3.
@@
The Marchenko–Pastur law with explicit estimates for the reminders (see e.g., Bourgade et al. (2022)), combined with the interlacing inequalities between the eigenvalues of spiked and unspiked models, as in Corollaries 9.8 and 9.13, can be used to show that
@@ 3.5. Implications and interpretations
Confidence intervals play two key roles in the analyzing signal strength. First, they measure uncertainty: the narrower the interval, the more precisely the signal is estimated.
Second, they assess the informativeness of estimated signals. If the lower bound starts at -$\infty$ or at the minimal admissible value of $\theta$, (which is $\sigma^{2}$ for the spiked covariance and zero for other models), the signal may be spurious and reflect pure noise. Alternatively, if the interval is bounded away from the minimal admissible value, but contains the identification threshold $\theta^{c}$, then a signal exists, but we cannot reject that its strength falls below the cutoff. When $\theta\leq\theta^{c}$ the sample estimates of the directions $\mathbf{u}$ and $\mathbf{v}$ are asymptotically orthogonal to their true counterparts (see, e.g., Paul (2007); Onatski (2012); Benaych-Georges and Nadakuditi (2012); Johnstone and Paul (2018); Bykhovskaya and Gorin (2025)), rendering the signal effectively non-informative.
@@ 4. Asymptotics through the Airy–Green function
The new asymptotics, which improves upon the Gaussian approximations (2.3), (2.9), (2.13), and (2.17), is based on a novel stochastic object we call the Airy–Green function. Its definition, along with the transition process $\mathcal{T}(\Theta)$ constructed from it, is presented in Section 4.1. A discussion of its nature is provided in Section 4.2. Theorem on the convergence of the eigenvalue distributions in four models towards this object is stated in Section 4.3.
@@ 4.1. Definition of $\mathcal{G}(w)$ and $\mathcal{T}(\Theta)$
Let $Y_{N}$ be an $N\times N$ matrix of i.i.d. $\mathcal{N}(0,\tfrac{2}{N})$ Gaussian random variables and let $\lambda_{1;N}\geq\lambda_{2;N}\geq\dots\geq\lambda_{N;N}$ be the eigenvalues of $\mathbf{B}=\frac{1}{2}\left(Y_{N}+Y_{N}^{\mathsf{T}}\right)$. Then in finite-dimensional distributions
@@
Similar asymptotic results hold for the other models we consider. All existing formulae for the finite-dimensional distributions of $\{\mathfrak{a}_{i}\}_{i=1}^{\infty}$ are quite complicated and do not provide explicit distribution function, see, e.g., Forrester (2010). Nevertheless, the distribution can be sampled and tabulated, see Bornemann (2009); Bykhovskaya et al. (2025). In particular, Table 2 lists quantiles of the Tracy–Widom distribution, which describes the law of $\mathfrak{a}_{1}$.
@@
(4.2) $\mathcal{G}(w)=\lim_{x\to-\infty}\left[\left(\sum_{j:\,\mathfrak{a}_{j}>x}\frac{\xi_{j}^{2}}{w-\mathfrak{a}_{j}}\right)-\frac{2}{\pi}\sqrt{-x}\right]\penalty 10000\ ,$
@@
Note that any fixed $w\in\mathbb{C}$ is almost surely not in $\{\mathfrak{a}_{j}\}$, hence, for such $w$ the convergence holds almost surely. Eq. (4.2) and Proposition 8.3 imply that $\mathcal{G}(w)$ changes monotonically from $\infty$ to -$\infty$ over the interval $\mathfrak{a}_{1},+\infty)$, allowing us to state the following key definition.
@@
Figure 2 and Table 1 show the simulated quantiles for the random variables $\mathcal{T}(\Theta)$ as functions of $\Theta$, or equivalently, confidence intervals for $\Theta$ as a function of $\mathcal{T}$. These results are based on $MC=10^{6}$ Monte Carlo simulations of the $\sqrt{N}\times\sqrt{N}$ top-left corners of $N\times N$ tridiagonal matrices of Dumitriu and Edelman (2002), with a perturbed $(1,1)$ matrix element and $N=10^{8}$; see Edelman and Persson (2005, Section 1.1) and Johnstone et al. (2021, Lemma 5.2) for justifications of this approach. The figure shows that the Gaussian approximation from Proposition 4.3 performs well for large $\Theta$, but deteriorates near $\Theta=0$.
@@ 4.2. Discussion of the definition
The term “Airy” in the name $\mathcal{G}(w)$ refers to the Airy point process, whose points $\mathfrak{a}_{i}$ appear in its definition. The term “Green” stands from the tradition in random matrix theory to refer to matrix elements of the resolvent $(zI-D)^{-1}$ of a symmetric matrix $D$ as the Green’s function. Via eigenvalue decomposition, the $(1,1)$ matrix element of $(zI-D)^{-1}$ is $\sum_{i}\frac{u_{1i}^{2}}{z-d_{i}}$, where $u_{1i}$ is the first coordinate of the $i$th normalized eigenvector of $D$ corresponding to the eigenvalue $d_{i}$, making it reminiscent of the sum in (4.2).
The term “transition” in $\mathcal{T}(\Theta)$ refers to its role in capturing the transition between subcritical $\theta<\theta^{c}$ and supercritical $\theta>\theta^{c}$ behavior in (2.3), (2.9), (2.13), (2.17). This phenomenon is commonly known as the BBP phase transition, following Baik et al. (2005).
@@ 4.3. Universal asymptotics for spiked models
(4.6) $\sqrt{N}(\lambda_{i}-\lambda(\theta_{i}))\xrightarrow[]{d}\mathcal{N}(0,V(\theta_{i})),\qquad 1\leq i\leq q-1,$
@@
(4.8) $\kappa_{1}=\frac{1}{2}\left[{\frac{[V^{\prime}(\theta^{c})]^{2}}{\lambda^{\prime\prime}(\theta^{c})}}\right]^{\frac{1}{3}},\qquad\kappa_{2}=\left[{\frac{[\lambda^{\prime\prime}(\theta^{c})]^{2}}{V^{\prime}(\theta^{c})}}\right]^{\frac{1}{3}}\penalty 10000\ .$
@@
While the distributional limit of $\lambda_{q+1},\dots,\lambda_{r}$ can be computed, it is of no use for the confidence intervals: the limits depend on $\tilde{\theta}$, but not on $\theta_{q+1},\dots,\theta_{r}$.
@@
If one sets $\theta_{i}=\theta^{c}+\varepsilon$, then Taylor expanding (4.6) (noting $V(\theta^{c})=\lambda^{\prime}(\theta^{c})=0$), we expect
@@
which is readily seen to be equivalent to (4.6). When $\theta_{i}$ is close to $\theta^{c}$, $\theta_{i}=\theta^{c}+N^{-1/3}\tilde{\theta}$, Taylor expanding $\lambda(\cdot)$ and $V(\cdot)$ near $\theta^{c}$, (4.9) turns into an equivalent form of (4.7):
@@
Since $\kappa_{3}/N=o(N^{-2/3})$, and therefore the choice of $\kappa_{3}$ does not affect the validity of the asymptotic formulas (4.9) and (4.10). These terms are introduced to improve the performance of the formulas for intermediate values of $N$, cf. Johnstone (2008); Ma (2012); Johnstone and Ma (2012), which emphasize the importance of $1/N$ corrections for the practical applicability. The reasoning behind our choice of $\kappa_{3}$ is as follows. First, we require continuity at $\theta^{c}$; hence, (4.9) and (4.10) use the same $\kappa_{3}/N$. Second, we leverage additional information available at $q=r=1$ and $\tilde{\theta}=-N^{1/3}\theta^{c}$ for the spiked Wigner, factor, and CCA models. (For the spiked covariance model, one instead takes $\tilde{\theta}=-N^{1/3}\gamma$ and adjusts the formula accordingly.) On one hand, combining (4.10) with the asymptotic approximation (4.4), we obtain
@@
On the other hand, $\tilde{\theta}=-N^{1/3}\theta^{c}$ corresponds to $\theta=0$ (or $\theta=1$ for the spiked covariance model with $\tilde{\theta}=-N^{1/3}\gamma$), meaning that in all four models, we are in the unspiked regime without a signal component. In this setting, the convergence of $\lambda_{1}$ to the Tracy–Widom distribution $\mathfrak{a}_{1}$ is well established, and $1/N$-order asymptotic corrections have been studied in Johnstone (2008); Ma (2012); Johnstone and Ma (2012). From these works, one can extract
[…]
@@
We expect that (4.9) also remains asymptotically valid on all mesoscropic scales, i.e., when $\theta_{i}=\theta^{c}+N^{-\alpha}\tilde{\theta}$, $0<\alpha<1/3$. We omit a detailed proof.
The proof of Theorem 4.6 in Section 9 begins with a rank-one perturbation equation, which expresses the eigenvalues in a signal plus noise model with $r$ spikes (the “target model”) as solutions to an algebraic equation involving a simpler model with r1r-1 spikes (the “base model”). Analyzing the asymptotic behavior of this equation leads to the following conclusion, stated informally below:
@@
We expect that our methods can be extended to handle the case of multiple ($k>1$) critical spikes, as well as the remaining largest eigenvalues $\lambda_{q+1},\lambda_{q+2},\dots$. The limiting behavior should be described by a higher-rank Airy point process, which we define recursively. The rank 0 process is the classical Airy point process $\{\mathfrak{a}_{j}\}$. The rank $1$ process $\{\mathfrak{a}_{j}^{(\Theta)}\}$ consists of all real solutions to the equation $\mathcal{G}(w)=-\Theta$; in particular, the largest point $\mathfrak{a}_{1}^{(\Theta)}$ coincides with $\mathcal{T}(\Theta)$ from Definition 4.2. We then iterate this construction: given the rank $k$ point process $\{\mathfrak{a}_{j}^{(\Theta_{1},\dots,\Theta_{k})}\}$ depending on $k$ real parameters $\Theta_{1},\dots,\Theta_{k}$, we define the rank $(k+1)$ process $\{\mathfrak{a}_{j}^{(\Theta_{1},\dots,\Theta_{k},\Theta_{k+1})}\}$ as the set of all real solutions to the equation
@@ 5. Empirical illustrations
We present three examples that illustrate the application of the procedure described in Section 3 to empirical data sets. In each case, the first step reveals a strong agreement between the histogram of eigenvalues and the corresponding theoretical curve—specifically, the Marchenko-Pastur law in the first two examples (factor models) and the Wachter law in the third (CCA). This suggests that the data aligns well with our modeling assumptions.
@@ 5.1. Industrial Production
Industrial production (IP) accounts for more than $10\%$ of the United States’ Gross Domestic Product (GDP), making it a significant component of total U.S. output. In this subsection we use data from Andreou et al. (2019), which investigates whether IP constitutes a dominant factor in U.S. economic activity. The data set contains quarterly IP growth rates across 117 sectors, spanning the period from $1977:Q1$ to $2011:Q4$. We de-mean the data and standardize each sector to have unit sample variance, thereby working with the sample correlation matrix of IP.
Figure 3 shows all eigenvalues of the standardized IP and highlights four of them (3.75, 4.26, 6.12, 30.053.75,\,4.26,\,6.12,\,30.05) that lie to the right of the theoretical Marchenko–Pastur upper edge, $\lambda_{+}=3.68$. The largest eigenvalue, $30.05$, stands out markedly and represents a strong “market” factor. The two smallest among the four, being close to the cutoff, may reflect spurious signals arising from noise. To assess their significance, we construct $95\%$ confidence intervals. These intervals are represented by vertical segments in Figure 3 and are also summarized in Table 5.
Notably, the interval for the fourth largest eigenvalue at $3.86$ differs substantially from what one would obtain using a Gaussian approximation (discussed in Section 2.1 and at the end of Section 2.3), highlighting the importance of our new procedure. This interval intersects the critical identification threshold $\theta^{c}=0.92$, indicating that we cannot reject the null hypothesis that it represents noise (or a non-informative signal). For the remaining eigenvalues, the null is rejected. The confidence interval for the largest eigenvalue is nearly identical under our method and the Gaussian approximation, whereas the differences grow as the eigenvalues and corresponding signal strengths decrease.
The discussion in the previous paragraph leads to the conclusion that the IP growth rate is driven by three factors: one strong factor and two weak factors. It is instructive to compare this result with the classical information criterion of Bai and Ng (2002), one of the most commonly used methods in applied research and one that is based on strong-signal asymptotics. Using $IC_{p_{2}}$ from Bai and Ng (2002, Section 5), we identify only the largest market factor. If we instead use $IC_{p_{1}}$ from the same article, which imposes a smaller penalty on the number of factors, we select the two largest factors. As a result, informative weak factors are missed by these procedures, underscoring the importance of accounting for weak signals and motivating our proposed methodology.
@@
Confidence Intervals
$\lambda_{i}$ $\theta_{i}$ Our method Gaussian method $\sharp 1$ Gaussian method $\sharp 2$
30.05 28.18 [26.48  30.07] [26.39  29.98] [26.44  30.03]
6.12 4.07 [3.41  4.94] [3.31  4.84] [3.37  4.90]
4.26 1.99 [1.41  2.65] [1.35  2.63] [1.44  2.70]
3.75 1.22 [0  2.02] [0.48  1.96] [0.93  1.99]
Table 5. IP: Confidence intervals based on different methods.